OpenStudy (anonymous):

Giving medal! A painter places an 11-ft ladder against a house. The base of the ladder is 3 ft from the house. How high on the house does the ladder reach? a. about 9.8 ft c. about 5.6 ft b. about 10.6 ft d. about 6. 1 ft

OpenStudy (lncognlto):

Do you know Pythagoras's theorem?

OpenStudy (anonymous):

nope

OpenStudy (lncognlto):

Well, here it is: Now C^2 = A^2 + B^2 (pythagoras) So 11^2 = 3^2 + B^2 Now we solve for B.

OpenStudy (anonymous):

how do we solve for b?

OpenStudy (lncognlto):

Well, we get it on its own, like this: 11^2 = 3^2 + B^2 therefore B^2 = 11^2 - 3^2 Now taking the square root of both sides, B = \[\sqrt{11^{2}-3^{2}}\]

OpenStudy (lncognlto):

Now using a calculator, B = ?

OpenStudy (anonymous):

i really dont get it im confused!

OpenStudy (lncognlto):

Pythagoras's theorem (a rule for triangles) states that the square of the length of the hypotenuse is equal to the square of the base plus the square of the height. So this is the ladder